Acetone, the solvent in many nail polish removers, has a density of 0.791 g/mL. What is the volume, in mL, of 74.7 g of acetone?
1
Expert's answer
2013-10-10T10:15:37-0400
V=m/r=74.7/0.791=94.4 mL